# YITP-SB-2025-20

### v1

@@ Abstract
We explore some aspects of the generalized Schur limit, defined in[Deb:2025ypl]. Based on several examples, we conjecture that the generalized Schur limit as a function of $\alpha$ solves a modular linear differential equation of fixed order, with coefficients depending on $\alpha$. We also observe in examples that for Argyres-Douglas theories of type $(A_{1},G)$ with $G=A_{n},D_{n}$, the generalized Schur limit for certain negative integer values of $\alpha$, coincides with the trace of higher powers of the quantum monodromy operator. This hints at a more general correspondence between the wall-crossing invariant traces on the Coulomb branch and the generalized Schur limit, which is related to the Higgs branch.
@@ 1 Introduction and summary
Four dimensional superconformal field theories (SCFTs) with eight supercharges have been a theoretical laboratory for studying the interplay between physics and mathematics. The superconformal index [Romelsberger:2005eg, Kinney:2005ej, Dolan:2008qi, Gadde:2011uv], which encodes information about the spectrum of the protected operators has been a major source of results relating to mathematical physics. One particular limit of the index is called the Schur index, which counts certain $1/4$-BPS operators, obeying certain shortening conditions on their quantum numbers [Gadde:2011ik, Gadde:2011uv, Rastelli:2014jja]. In particular, to a 4d SCFT one can associate a 2d vertex operator algebra (VOA) and the Schur index equals the vaccum character of the corresponding VOA [Beem:2013sza]. This limit has several interesting modular properties. The Schur index is conjectured to be a vector valued modular form and solves a modular linear differential equation (MLDE) [Arakawa:2016hkg, Beem:2017ooy]. Here we will examine properties of a limit closely related to the Schur index.
Recently in [Deb:2025ypl], a double scaling limit of the superconformal index, called there as generalized Schur limit and denoted by $\hat{\mathcal{Z}}(q,\alpha)$, was defined. This limit is parametrized by a parameter $\alpha\geq 0$ and for a subset of theories, at special values of this parameter, the limit reduces to the Schur index of theories related to the initial theory by renormalization group flows. In certain cases, this limit provides a very simple contour integral expression for the Schur index indices of non-Lagrangian SCFTs.  [^1] Using this limit [Pan:2025vyu] computed a closed form expression for partially flavored Schur index of rank-one $E_{6}$ and $E_{7}$ Minahan-Nemeschansky theories.. This limit is well defined for any $\alpha\geq 0$, but it may or may not give rise to a integer coefficient $q$-series. In this paper we explore the range $\alpha<0$. To analyze this range, we evaluate the index as a $q$-series with $\alpha$ dependent coefficients and use the expression to analytically continue to negative $\alpha$. In this range, for certain special values of $\alpha$, we find that the index equals the vacuum character of vertex operator algebras, which show up in the trace of higher powers of wall-crossing invariant quantum monodromy operator222[Cordova:2015nma] used the operator $\mathcal{O}(q)$ called the Kontsevich-Soibelman operator [Kontsevich:2008fj], which is related to the quantum monodromy as $\mathcal{O}(q)=M(q)^{-1}$. $M(q)$ [Kim:2024dxu]. We test this idea in the case of Argyres-Douglas theories of type $(A_{1},G)$, where $G=A_{n},D_{n}$ . In the examples, we find that the generalized Schur partition function equals the trace of a higher power $M(q)$  [^3] A similar limit was considered in [Cecotti:2015lab] and its relation to the trace of $M(q)$ was discussed.. To be more explicit
@@
The two sides of equation (1.1) seem to connect two different quantities. The moduli space of vacua of 4d $\mathcal{N}=2$ SCFTs has two distinct branches: the Higgs branch and the Coulomb branch. The left hand side is a quantity related to the Higgs branch. This is because the Schur operators contain Higgs branch operators as a subset and do not contain Coulomb branch operators. The right hand side is a wall crossing invariant defined using the data of BPS states on the Coulomb branch of the theory. [Cordova:2015nma] observed that in a large class of theories, the Schur index is to equal $(q;q)^{2\mathfrak{r}}\mbox{Tr}M(q)^{-1}$, thus hinting at a relation between the two seemingly distinct branches (see also [Cordova:2016uwk, Cordova:2017ohl, Cordova:2017mhb, Cecotti:2010fi, Kaidi:2022sng, Buican:2024jjl, Fredrickson:2017yka, Pan:2024hcz, Deb:2025cqr, Li:2025nhc] ). Therefore the equality (1.1) seems to hint at a generalization of the correspondence of [Cordova:2015nma] . As a naive check of equation (1.1), note that the conjectured central charge of the 4d SCFT obtained in the generalized Schur limit of [Deb:2025ypl] matches that of the central charge conjectured in [Kim:2024dxu] with the identification $\alpha=-N$
@@
In order to check equation (1.1), we need to compute the generalized Schur limit for negative values of $\alpha$. We will restrict ourselves to $(A_{1},G)$ type theories, with $G=A_{n},D_{n}$ , whose index is available as a contour integral obtained from deforming a $\mathcal{N}=1$ Lagrangian theory [Maruyoshi:2016tqk, Maruyoshi:2016aim, Agarwal:2016pjo]. The generalized Schur limit for these Argyres-Douglas theories turns out to be equal to the generalized Schur limit of some $\mathcal{N}=2$ SQCD by a rescaling of $\alpha$ by some number $\xi$ i.e.
@@
This brings us to another approach of using the MLDE. As mentioned above, the Schur index solves an MLDE. A finite number of numerical coefficients are sufficient to define an MLDE and therefore the Schur index can be computed upto arbitrary high orders only with this finite amount of data. In [Deb:2025ypl], it was noted that for all values of $\alpha$ where the generalized Schur limit could be computed explicitly as $q$-series expansion, the $q$-series thus obtained solved a modular linear differential equation (MLDE) of the same order. We find this to be true in all the examples we have considered. Here, we will show that this is true for the case of $SU(2)$ with four flavors, in a $q$-series expansion upto some order. For higher rank cases we check this statement for positive $\alpha$ and find it to be true in whichever value this check could be performed. Based on this, we conjecture the following
@@
Assuming this conjecture, the generalized Schur limit of an integral reduces to finding a modular linear differential operator (MLDO) that annihilates the Schur index for all values of $\alpha$. To determine this MLDO, we will first plot the coefficients of the respective MLDO as a function of $\alpha$ and perform a fit to obtain the MLDO as a function of $\alpha$. Solving this MLDE gives the generalized Schur index as a function of $\alpha$. We use the expression thus obtained to explore the range $\alpha<0$. In most general case, the cofficients of the $q$-series and MLDE can be arbitrary functions of $\alpha$, but in all the examples we consider here, we find that the coefficients of $q$-series are rational functions of $\alpha$ and the coefficients of MLDE are polynomials in $\alpha$. As the rank of the theory increases, the order of the MLDE is also expected to increase [Deb:2025cqr]. Therefore, for checks upto very low orders in $q$, the method of directly fitting the $q$-series coefficients is sometimes better and we use this technique to perform checks by refining the index by the flavor fugacities for higher rank theories.
@@
Even though we have verified the above claims in a limited number of examples, the results seem to indicate a more general structure. A concrete mathematical problem is to show rigorously that at least in the case of Lagrangian theories, where the generalized Schur limit can be taken, the corresponding MLDE is of fixed order with coefficients depending on $\alpha$. Currently, this limit can be conviently evaluated for theories which have a Lagragian description or atleast a contour integral for the index using a Lagrangian that flows to them in the infrared. It would be interesting to show this for other non Lagrangian theories. In these computations we encoutered integer coefficient $q$-series, which were not able to identify with known VOAs. It would be interesting to determine whether these correspond to a genuine VOA. With respect to MLDE computations, we mainly discuss the unflavored case but it would be interesting to check the properties of flavored MLDEs for the generalized Schur limit with all the flavors fugacities [Zheng:2022zkm, Pan:2023jjw].
@@
We only do the examples related to $(A_{1},G)$ type Argyes-Douglas theories, but the numerical fit technique for determining MLDEs or the coefficient of $q$-series is quite general and can be easily applied to various examples. Using fixed order MLDE to bootstrap the characters of 2d RCFTs is an old subject [Mathur:1988rx, Mukhi:2019xjy, Mukhi:2019cpu, Mukhi:2020gnj] and similar in spirit to computations in this paper. We will discuss the rank-one DC in detail in Section 3, but a natural question to ask is what happens for higher rank DC series. Interestingly, the rank-two series solves a fourth-order MLDE parametrized by a single parameter. One also obtains integer $q$-series for certain rational values of this parameter, as well as the rank-two analouges of $\mathfrak{a}_{\frac{1}{2}}$ and $\mathfrak{e}_{7\frac{1}{2}}$, resembling the structure in the rank one case. But this does not coincide with the generalized Schur limit of the $USp(4)$ Lagrangian theory in that series. This seems to indicate that there might be some interesting integral formula for the higher rank DC series, but it may or may not be some limit of the full superconformal index. It would be interesting to explore this in future.
@@
Note: While we were in the final stages of the project, we were informed about another group studying similar topics with some overlap. This paper is being submitted in coordination with their paper [toappcoord]. We thank its authors for their cooperation.
@@ 2.1 Generalized Schur limit
The superconformal index of a 4d $\mathcal{N}=2$ SCFTs is given by [Romelsberger:2005eg, Kinney:2005ej, Dolan:2008qi, Gadde:2011uv]
@@
where $p,q$ and $t$ are superconformal fugacities and $u_{i}$ are the flavour fugacities. $R$ is the Cartan generator of the $SU(2)_{R}$ R-symmetry, $r$ is the $U(1)_{r}$ R-charge, and $j_{i}$ are the Cartan generators of the $SU(2)\times SU(2)$ isometry of $\mathbb{S}^{3}$. The limit $t=q$ is called the Schur index and in this limit, the index is independent of $p$ [Gadde:2011ik, Gadde:2011uv, Rastelli:2014jja]
@@
To every 4d $\mathcal{N}=2$ SCFT, one can associate a vertex operator algebra (VOA) [Beem:2013sza]. As a vector space, the corresponding VOA captures the Schur operators and the vacuum character of the VOA equals the Schur index of the SCFT. The Higgs branch operators of the theory are a subset of the Schur operators. In fact, the Higgs branch reconstruction conjecture asserts that the Higgs branch can be recovered from the VOA by the associated variety construction [Arakawa:2016hkg, Beem:2017ooy]. Assuming the Higgs branch reconstruction conjecture, it can be shown that the VOA vacuum character must obey a monic modular linear differential equation (MLDE). Therefore Schur indices are expected to be solutions to MDLEs.
@@
$\mathcal{I}_{\text{hyp}}$ and $\mathcal{I}_{\text{vec}}$ denote the indices for the vector and hypermultiplet, which can be written as the plethystic exponential666 $\text{PE}\left[f(x_{1},x_{2},\dots,x_{n})\right]:=e^{\sum_{k=1}^{\infty}\frac{1}{k}f(x_{1}^{k},x_{2}^{k},\dots,x_{n}^{k})}$ of single particle indices as follows
@@
$\mathcal{I}_{\text{vec}}(\mathbf{a},q)$ $=\text{PE}\left[\frac{2q}{1-q}\chi^{\mathfrak{g}}_{\text{adj}}(\mathbf{a})\right]~.$ (2.8)
$\chi^{\mathfrak{g}}_{R}(\mathbf{a})$ denotes the character of representation $R$ of the half-hypermultiplets in the gauge algebra $\mathfrak{g}$. Note that in this limit, the Haar measure $\Delta(\mathbf{a})$ is raised to the power $\alpha$. It has been observed in a variety of examples that, for certain positive values of $\alpha$, this limit equals the Schur index of theories related to this theory by RG flows. It was also conjectured in [Deb:2025ypl] that a necessary condition for the partition function of two theories $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ to be related to each other as follows
@@
$c_{\text{4d}}{(2)}=c_{\text{4d}}^{(1)}\alpha+\frac{1}{6}(1-\alpha)\mathfrak{r}\,,\;\;\;\Delta^{(2)}_{i}=(\Delta^{(1)}_{i}-1)\alpha+1~.$ (2.10)
@@ 2.2 Trace of quantum monodromy
The moduli space of vacua of 4d $\mathcal{N}=2$ SCFTs have two distinct branches, the Higgs branch and the Coulomb branch. These have two seemigly different mathemtical invariants associated to them. The Coulomb branch is described the special Kähler geometry [Seiberg:1994rs, Seiberg:1994aj] and as discussed above, the Higgs branch equals the associated variety of the VOA. Apriori, the Higgs and Coulomb branch seem unrelated. In [Cordova:2015nma], a non trivial relation between the Higgs and Coulomb branch was observed where the trace of the inverse of quantum monodromy $M(q)$, was shown to be equal to the Schur index of the theory
@@
The flavor symmetries $G_{F}$ and flavor lattice vectors of $(A_{1},G)$ theories are collected in Table 1. We refer the reader to [Kim:2024dxu, Cordova:2015nma, Cecotti:2015lab] for more details on computing these traces.
@@
$(A_{1},A_{2N+1})$
$SU(2)$, $N=1$
$U(1)$, $N>1$
$\gamma_{f}=\sum_{i=0}^{n}(-1)^{i}\gamma_{2i+1}$ $\mbox{Tr}X_{\gamma_{f}}=\begin{cases}a^{2},~N=1\\ (-1)^{N+1}a,~N>1\end{cases}$
@@
$(A_{1},D_{2N+2})$
$SU(3)$, $N=1$
$SU(2)\times U(1)$, $N>1$
$\gamma_{f_{1}}=\gamma_{2n+2}-\gamma_{2n+1}$
$\gamma_{f_{2}}=\sum_{i=0}^{n}(-1)^{i}\gamma_{2i+1}$
$\mbox{Tr}X_{\gamma_{f_{1}}}=b^{2}$
$\mbox{Tr}X_{\gamma_{f_{1}}}=\begin{cases}ab,~N~\text{odd}\\ -\frac{a}{b},~N~\text{even}\end{cases}$

### v2

@@ Abstract
We explore some aspects of the generalized Schur limit, defined in[21]. Based on several examples, we conjecture that the generalized Schur limit as a function of $\alpha$ solves a modular linear differential equation of fixed order, with coefficients depending on $\alpha$. We also observe in examples that for Argyres-Douglas theories of type $(A_{1},G)$ with $G=A_{n},D_{n}$, the generalized Schur limit for certain negative integer values of $\alpha$, coincides with the trace of higher powers of the quantum monodromy operator. This hints at a more general correspondence between the wall-crossing invariant traces on the Coulomb branch and the generalized Schur limit, which is related to the Higgs branch.
@@ 1 Introduction and summary
Four dimensional superconformal field theories (SCFTs) with eight supercharges have been a theoretical laboratory for studying the interplay between physics and mathematics. The superconformal index [44, 29, 22, 26], which encodes information about the spectrum of the protected operators has been a major source of results relating to mathematical physics. One particular limit of the index is called the Schur index, which counts certain $1/4$-BPS operators, obeying certain shortening conditions on their quantum numbers [25, 26, 43]. In particular, to a 4d SCFT one can associate a 2d vertex operator algebra (VOA) and the Schur index equals the vaccum character of the corresponding VOA [5]. This limit has several interesting modular properties. The Schur index is conjectured to be a vector valued modular form and solves a modular linear differential equation (MLDE) [2, 6]. Here we will examine properties of a limit closely related to the Schur index.
Recently in [21], a double scaling limit of the superconformal index, called there as generalized Schur limit and denoted by $\hat{\mathcal{Z}}(q,\alpha)$, was defined. This limit is parametrized by a parameter $\alpha\geq 0$ and for a subset of theories, at special values of this parameter, the limit reduces to the Schur index of theories related to the initial theory by renormalization group flows. In certain cases, this limit provides a very simple contour integral expression for the Schur index indices of non-Lagrangian SCFTs.  [^1] Using this limit [42] computed a closed form expression for partially flavored Schur index of rank-one $E_{6}$ and $E_{7}$ Minahan-Nemeschansky theories.. This limit is well defined for any $\alpha\geq 0$, but it may or may not give rise to a integer coefficient $q$-series. In this paper we explore the range $\alpha<0$. To analyze this range, we evaluate the index as a $q$-series with $\alpha$ dependent coefficients and use the expression to analytically continue to negative $\alpha$. In this range, for certain special values of $\alpha$, we find that the index equals the vacuum character of vertex operator algebras, which show up in the trace of higher powers of wall-crossing invariant quantum monodromy operator222[18] used the operator $\mathcal{O}(q)$ called the Kontsevich-Soibelman operator [30], which is related to the quantum monodromy as $\mathcal{O}(q)=M(q)^{-1}$. $M(q)$ [28]. We test this idea in the case of Argyres-Douglas theories of type $(A_{1},G)$, where $G=A_{n},D_{n}$ . In the examples, we find that the generalized Schur partition function equals the trace of a higher power $M(q)$  [^3] A similar limit was considered in [13] and its relation to the trace of $M(q)$ was discussed.. To be more explicit
@@
The two sides of equation (1.1) seem to connect two different quantities. The moduli space of vacua of 4d $\mathcal{N}=2$ SCFTs has two distinct branches: the Higgs branch and the Coulomb branch. The left hand side is a quantity related to the Higgs branch. This is because the Schur operators contain Higgs branch operators as a subset and do not contain Coulomb branch operators. The right hand side is a wall crossing invariant defined using the data of BPS states on the Coulomb branch of the theory. [18] observed that in a large class of theories, the Schur index is to equal $(q;q)^{2\mathfrak{r}}\mbox{Tr}M(q)^{-1}$, thus hinting at a relation between the two seemingly distinct branches (see also [15, 16, 17, 12, 27, 11, 24, 41, 20, 31, 8] ). Therefore the equality (1.1) seems to hint at a generalization of the correspondence of [18] . As a naive check of equation (1.1), note that the conjectured central charge of the 4d SCFT obtained in the generalized Schur limit of [21] matches that of the central charge conjectured in [28] with the identification $\alpha=-N$
@@
In order to check equation (1.1), we need to compute the generalized Schur limit for negative values of $\alpha$. We will restrict ourselves to $(A_{1},G)$ type theories, with $G=A_{n},D_{n}$ , whose index is available as a contour integral obtained from deforming a $\mathcal{N}=1$ Lagrangian theory [33, 32, 1]  [^5] The indices of such theories were also obtained using other techniques. See for example [9, 10]. . The generalized Schur limit for these Argyres-Douglas theories turns out to be equal to the generalized Schur limit of some $\mathcal{N}=2$ SQCD by a rescaling of $\alpha$ by some number $\xi$ i.e.
@@
This brings us to another approach of using the MLDE. As mentioned above, the Schur index solves an MLDE. A finite number of numerical coefficients are sufficient to define an MLDE and therefore the Schur index can be computed upto arbitrary high orders only with this finite amount of data. In [21], it was noted that for all values of $\alpha$ where the generalized Schur limit could be computed explicitly as $q$-series expansion, the $q$-series thus obtained solved a modular linear differential equation (MLDE) of the same order. We find this to be true in all the examples we have considered. Here, we will show that this is true for the case of $SU(2)$ with four flavors, in a $q$-series expansion upto some order. For higher rank cases we check this statement for positive $\alpha$ and find it to be true in whichever value this check could be performed. Based on this, we conjecture the following
@@
Assuming this conjecture, the generalized Schur limit of an integral reduces to finding a modular linear differential operator (MLDO) that annihilates the Schur index for all values of $\alpha$. To determine this MLDO, we will first plot the coefficients of the respective MLDO as a function of $\alpha$ and perform a fit to obtain the MLDO as a function of $\alpha$. Solving this MLDE gives the generalized Schur index as a function of $\alpha$. We use the expression thus obtained to explore the range $\alpha<0$. In most general case, the cofficients of the $q$-series and MLDE can be arbitrary functions of $\alpha$, but in all the examples we consider here, we find that the coefficients of $q$-series are rational functions of $\alpha$ and the coefficients of MLDE are polynomials in $\alpha$. As the rank of the theory increases, the order of the MLDE is also expected to increase [20]. Therefore, for checks upto very low orders in $q$, the method of directly fitting the $q$-series coefficients is sometimes better and we use this technique to perform checks by refining the index by the flavor fugacities for higher rank theories.
@@
Even though we have verified the above claims in a limited number of examples, the results seem to indicate a more general structure. A concrete mathematical problem is to show rigorously that at least in the case of Lagrangian theories, where the generalized Schur limit can be taken, the corresponding MLDE is of fixed order with coefficients depending on $\alpha$. Currently, this limit can be conviently evaluated for theories which have a Lagragian description or atleast a contour integral for the index using a Lagrangian that flows to them in the infrared. It would be interesting to show this for other non Lagrangian theories. In these computations we encoutered integer coefficient $q$-series, which were not able to identify with known VOAs. It would be interesting to determine whether these correspond to a genuine VOA. With respect to MLDE computations, we mainly discuss the unflavored case but it would be interesting to check the properties of flavored MLDEs for the generalized Schur limit with all the flavors fugacities [47, 40].
@@
We only do the examples related to $(A_{1},G)$ type Argyes-Douglas theories, but the numerical fit technique for determining MLDEs or the coefficient of $q$-series is quite general and can be easily applied to various examples. Using fixed order MLDE to bootstrap the characters of 2d RCFTs is an old subject [34, 39, 37, 38] and similar in spirit to computations in this paper. We will discuss the rank-one DC in detail in Section 3, but a natural question to ask is what happens for higher rank DC series. Interestingly, the rank-two series solves a fourth-order MLDE parametrized by a single parameter. One also obtains integer $q$-series for certain rational values of this parameter, as well as the rank-two analouges of $\mathfrak{a}_{\frac{1}{2}}$ and $\mathfrak{e}_{7\frac{1}{2}}$, resembling the structure in the rank one case. But this does not coincide with the generalized Schur limit of the $USp(4)$ Lagrangian theory in that series. This seems to indicate that there might be some interesting integral formula for the higher rank DC series, but it may or may not be some limit of the full superconformal index. It would be interesting to explore this in future.
@@
Note: While we were in the final stages of the project, we were informed about another group studying similar topics with some overlap. This paper is being submitted in coordination with their paper [14]. We thank its authors for their cooperation.
@@ 2.1 Generalized Schur limit
The superconformal index of a 4d $\mathcal{N}=2$ SCFTs is given by [44, 29, 22, 26]
@@
where $p,q$ and $t$ are superconformal fugacities and $u_{i}$ are the flavour fugacities. $R$ is the Cartan generator of the $SU(2)_{R}$ R-symmetry, $r$ is the $U(1)_{r}$ R-charge, and $j_{i}$ are the Cartan generators of the $SU(2)\times SU(2)$ isometry of $\mathbb{S}^{3}$. The limit $t=q$ is called the Schur index and in this limit, the index is independent of $p$ [25, 26, 43]
@@
To every 4d $\mathcal{N}=2$ SCFT, one can associate a vertex operator algebra (VOA) [5]. As a vector space, the corresponding VOA captures the Schur operators and the vacuum character of the VOA equals the Schur index of the SCFT. The Higgs branch operators of the theory are a subset of the Schur operators. In fact, the Higgs branch reconstruction conjecture asserts that the Higgs branch can be recovered from the VOA by the associated variety construction [2, 6]. Assuming the Higgs branch reconstruction conjecture, it can be shown that the VOA vacuum character must obey a monic modular linear differential equation (MLDE). Therefore Schur indices are expected to be solutions to MDLEs.
@@
$\mathcal{I}_{\text{hyp}}$ and $\mathcal{I}_{\text{vec}}$ denote the indices for the vector and hypermultiplet, which can be written as the plethystic exponential777 $\text{PE}\left[f(x_{1},x_{2},\dots,x_{n})\right]:=e^{\sum_{k=1}^{\infty}\frac{1}{k}f(x_{1}^{k},x_{2}^{k},\dots,x_{n}^{k})}$ of single particle indices as follows
@@
$\mathcal{I}_{\text{vec}}(\mathbf{a},q)$ $=\text{PE}\left[\frac{-2q}{1-q}\chi^{\mathfrak{g}}_{\text{adj}}(\mathbf{a})\right]~.$ (2.8)
$\chi^{\mathfrak{g}}_{\mathcal{R}}(\mathbf{a})$ denotes the character of representation $\mathcal{R}$ of the half-hypermultiplets in the gauge algebra $\mathfrak{g}$. Note that in this limit, the Haar measure $\Delta(\mathbf{a})$ is raised to the power $\alpha$. It has been observed in a variety of examples that, for certain positive values of $\alpha$, this limit equals the Schur index of theories related to this theory by RG flows. It was also conjectured in [21] that a necessary condition for the partition function of two theories $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ to be related to each other as follows
@@
$c_{\text{4d}}^{(2)}=c_{\text{4d}}^{(1)}\alpha+\frac{1}{6}(1-\alpha)\mathfrak{r}\,,\;\;\;\Delta^{(2)}_{i}=(\Delta^{(1)}_{i}-1)\alpha+1~.$ (2.10)
@@ 2.2 Trace of quantum monodromy
The moduli space of vacua of 4d $\mathcal{N}=2$ SCFTs have two distinct branches, the Higgs branch and the Coulomb branch. These have two seemigly different mathemtical invariants associated to them. The Coulomb branch is described the special Kähler geometry [45, 46] and as discussed above, the Higgs branch equals the associated variety of the VOA. Apriori, the Higgs and Coulomb branch seem unrelated. In [18], a non trivial relation between the Higgs and Coulomb branch was observed where the trace of the inverse of quantum monodromy $M(q)$, was shown to be equal to the Schur index of the theory
@@
The flavor symmetries $G_{F}$ and flavor lattice vectors of $(A_{1},G)$ theories are collected in Table 1. We refer the reader to [28, 18, 13] for more details on computing these traces.
@@
$(A_{1},A_{2N+1})$ $SU(2)$, $N=1$ $U(1)$, $N>1$ $\gamma_{f}=\sum_{i=0}^{N}(-1)^{i}\gamma_{2i+1}$ $\mbox{Tr}X_{\gamma_{f}}=\begin{cases}a^{2},~N=1\\ (-1)^{N+1}a,~N>1\end{cases}$
@@
$(A_{1},D_{2N+2})$ $SU(3)$, $N=1$ $SU(2)\times U(1)$, $N>1$ $\gamma_{f_{1}}=\gamma_{2N+2}-\gamma_{2N+1}$ $\gamma_{f_{2}}=\sum_{i=0}^{N}(-1)^{i}\gamma_{2i+1}$ $\mbox{Tr}X_{\gamma_{f_{1}}}=b^{2}$ $\mbox{Tr}X_{\gamma_{f_{2}}}=\begin{cases}ab,~N~\text{odd}\\ -\frac{a}{b},~N~\text{even}\end{cases}$